LOG-IN
Displaying report 1-1 of 1.
CAL (General)
Print this report.
chihiro.kozakai - 09:36, Monday 05 November 2018 (6848)Get code to link to this report
Termination of the calibration line injection test by PCal

Successfully finished the calibration line injection test.

I stopped the calibration line injection test (15 Hz with full range) in this morning. The laser modulation control was still alive from 11/3 evening ! Therefore, continuous operation of 38 hours (from 11/3 5pm to 11/5 8am) is achieved in this weekend.

The averaged spectrum is attached. The representative values are:

Noise floor
Tx1 -79 dBV
OFS1 -82 dBV
ToAOM1 -104 dBV
Tx2 -85 dBV
OFS2 -81 dBV
ToAOM2 -106 dBV

Chihiro

Images attached to this report
6848_20181105013327_16.png